A project has an initial outlay of $2,396. The project will generate annual cash flows of $593 over the 4-year life of the project and terminal cash flows of $285 in the last year of the project. If the required rate of return on the project is 11%, what is the net present value (NPV) of the project?
Answer #1

Calculation of NPV

Answer: NPV = -$368.55

NPV = Present Value of cash inflows – Initial Investment

Calculation of Present value of cash inflows for project

Year

Cash Flow

Present Value Factor @ 11%

Present Value of cash flow

(I)

(II)

(III)

(II) * (III)

1

$593

0.9009

$534.23

2

$593

0.8116

$481.28

3

$593

0.7312

$433.60

4

$593

0.6587

$390.61

4

$285

0.6587

$187.73

Present Value of the Cash flows inflows

$2,027.45

Initial Investment = $2,396/- (provided in the question)

NPV = Present Value of cash inflows – Initial Investment

NPV = $2,027.45 - $2,396

            = -368.55

Calculation of Discounting Factor (Present Value Factor)

Discount Factor = 1/ (1+R) N

R = Discount Rate (i.e. = 11%)

N = No of years

E.g. for year 2 Discount Factor = 1/ (1.11)2

                                                                = 1/ (1.11) (1.11)

                                                = 0.8116